OpenStudy (anonymous):

Please. Solve the following system of equations. 3x - 2y = 6 6x - 4y = 12 Choices: (0, 0) (6, 12) Infinitely Many Solutions No Solutions

OpenStudy (jdoe0001):

have you solved it yet?

OpenStudy (anonymous):

I think it's no solution

OpenStudy (anonymous):

Haven't solved yet, no

OpenStudy (anonymous):

I solved it and its no solution

OpenStudy (anonymous):

Thank you

OpenStudy (jdoe0001):

$$ (-6x+4x=-12)*\boldsymbol{(-2)}\\ 6x-46=12 $$

OpenStudy (jdoe0001):

shoot, so many typos :|

OpenStudy (jdoe0001):

$$ (-3x+2x=-6)*\boldsymbol{(-2)}\\ 6x-4=12 $$ which is what I meant, would give you an incosistent system

OpenStudy (anonymous):

ok that makes sense

OpenStudy (jdoe0001):

$$ (3x-2y=6)*\boldsymbol{(-2)}\\ 6x-4y=12 $$, sorry for the typos :(

OpenStudy (anonymous):

Same line. It is not an inconsistent system (parallel lines). It is a dependent system having infinitely many solutions. This becomes evident after multiplying both sides of equation 1 by "2". You end up with the same equation. That's all that needs to be done.

OpenStudy (anonymous):

Good luck to you in all of your studies and thx for the recognition! @marylou004

OpenStudy (anonymous):

-2(3x-2y=6) =-6x+4y=-12. (I used the elimination) 6x-4y=12 It's no solution cause you Need to cancel